# WZW terms without anomalies: generalised symmetries in chiral Lagrangians

### v1

@@ 1 Introduction
Fast forward half a century, and symmetry continues to offer rich and surprising insights into the dynamics of QFTs. A key development in the last decade has been the discovery of generalised notions of symmetry Gaiotto:2014kfa , beyond the action of groups on local operators. One example is higher $p$-form symmetries, in which abelian groups act not on local operators but on extended objects. From the viewpoint of current algebra, the corresponding symmetry currents for $p$-form symmetries (in the case of continuous symmetry) are not vectors $j^{(1)}_{\mu}$ (better, 1-forms) as they are for ‘ordinary’ 00-form symmetries, but tensors $j^{(p+1)}_{\mu...\sigma}$ with more than one index (specifically, $(p+1)$-forms). Returning to QCD and its discrete symmetries, $\text{SU}(n)$ gauge theory possesses such a $\mathbb{Z}_{n}$ valued 1-form symmetry, which was shown by Gaiotto, Kapustin, Komargodski and Seiberg to have a mixed anomaly with parity when the QCD theta angle equals $\pi$ Gaiotto:2017yup . This subtle anomaly in QCD, discovered nearly 50 years after the ABJ anomaly, means that QCD at $\theta=\pi$ cannot be trivial in the deep infrared (IR).
It has also been recently understood that there can be a non-trivial current algebra between higher-form symmetries of different degree. The simplest example is known as a 2-group symmetry structure, wherein 0-form and 1-form symmetries ‘mix’, as first detected in quantum field theories by Sharpe Sharpe:2015mja . In the case where both 0-form and 1-form symmetries are continuous, this can be captured, following the tradition of current algebra, by a Ward identity of the form Cordova:2020tij :
@@ 2 Mixed WZW term on $\mathrm{SU}(N)\times S^{2}$
(Co)homological Davighi:2018inx (or (co)bordism-based Lee:2020ojw ; Yonekura:2020upo ) classifications of Wess–Zumino–Witten (WZW) terms Wess:1971yu ; Witten:1983tw tell us there is a mixed WZW interaction coming from the existence of a $G$-invariant closed 5-form involving pions on both the $\text{SU}(N)$ and $S^{2}$ factors, namely Davighi:2024zip
@@
where $\mathrm{Vol}_{S^{2}}$ is the volume form on the $S^{2}$. To define this term it is easier to start with a homological description (we briefly discuss the refinement via bordism afterwards). Given there are no homologically non-trivial 4-cycles in the target space, since $H_{4}(\text{SU}(N)\times S^{2})=0$, any 4-cycle in the target space obtained by pushing forward $\Sigma_{4}$ (more precisely, pushing forward a cycle in the fundamental class $[\Sigma_{4}]$) can be realised as the boundary of a 5-cochain $X_{5}$, on which we can integrate $\omega$ to obtain the exponentiated action à la Witten Witten:1983tw
@@
$L=\frac{in\epsilon^{\mu\nu\rho\sigma}}{48\pi^{2}f_{D}^{2}f_{\pi}^{3}}f_{abc}\epsilon_{ij}\pi_{a}\partial_{\mu}\pi_{b}\partial_{\nu}\pi_{c}\partial_{\rho}\chi_{i}\partial_{\sigma}\chi_{j}+\mathcal{O}(\pi^{4}\chi^{2},\pi^{3}\chi^{3})\,,$ (7)
@@
2.1 Remarks on bordism vs homology
@@
Roughly speaking, bordism tells us whether manifolds equipped with certain structures (gauge bundles, spin structures, metrics) can be written as boundaries of manifolds in one dimension higher, with all structures smoothly extended thereto. In short, bordism deals directly with manifolds themselves (rather than passing to cochains as in the homological description above). The UV theory we have in mind will feature fermions, so the appropriate bordism theory is (reduced) spin-bordism.
@@
$\tilde{\Omega}_{4}^{\mathrm{Spin}}\left(\text{SU}(N)\times S^{2}\right)\cong\mathbb{Z}_{2}\,.$ (8)
This means there is a class of spin 4-manifolds $[X_{4}]$, together with maps $\sigma$ to $\text{SU}(N)\times S^{2}$, that cannot be extended to spin 5-manifolds (also equipped with maps to $\text{SU}(N)\times S^{2}$); this equivalence class is the generator of the bordism group (8). The existence of manifolds that are not boundaries is an obstruction to a Witten-like definition (3) of the WZW term when evaluated on such manifolds. This obstruction is not seen using homology which, amongst other differences, is not sensitive to spin structures and whether or not they can be extended also.
We can nonetheless proceed in attempting to define our WZW term on a general manifold as follows. That the bordism group is $\mathbb{Z}_{2}$ means that the union of two generators is a boundary of some 5-manifold $X_{5}$ with the map $\sigma$ extended. The exponentiated action evaluated on this union $X_{4}\sqcup X_{4}$ must be the square of the action evaluated on the original $X_{4}$, by locality. Therefore, we can define the exponentiated WZW action for this union as
@@
There are then two possibilities for the WZW term for the generator of the bordism group, differing by a sign. This sign can be fixed by anomaly matching: the solution with a minus sign must be chosen if the symmetry group $\text{SU}(2)$ suffers from the mod 2 global anomaly discovered by Witten Witten:1982fp in the UV. For more detail on matching global anomalies with WZW terms, see Refs. Yonekura:2020upo ; Lee:2020ojw .
@@ 2.2 An anomaly-matching puzzle
$L\sim\frac{1}{f_{\pi}f_{D}^{2}}\epsilon^{\mu\nu\rho\sigma}\pi_{0}F_{\mu\nu}\partial_{\rho}\chi_{1}\partial_{\sigma}\chi_{2}\,,$ (10)
@@
$L\sim\frac{1}{f_{\pi}}\epsilon^{\mu\nu\rho\sigma}\pi_{0}F_{\mu\nu}(F_{D})_{\rho\sigma}\,.$ (11)
@@ 2.3 Infrared 2-group symmetry
$iS[\Sigma_{4}=\partial X_{5}]=\int_{X_{5}}\frac{-i}{8\pi^{2}}\left[\mathrm{CS}(A_{L})-\mathrm{CS}(A_{R})\right]\wedge\mathrm{Vol}_{S^{2}}$ (15)
@@
Mathematically, this (together with 16) defines the infinitesimal version of a continuous 2-group gauge transformation, [^6] We will not actually define what a 2-group is in this paper. For an introduction to the mathematical notion of a 2-group, and how these structures appear in QFT, see e.g. §2 of Davighi:2023luh . with the locally-defined differential forms $(A_{L},A_{R},B)$ forming a 2-connection. Given the 0-form flavour symmetry $\text{SU}(N)^{2}$ and the 1-form winding number symmetry valued in $\mathrm{U}(1)^{[1]}$, possible 2-group structures intertwining the two are classified (subject to certain simplifying assumptions that apply here) by a topological invariant called the Postnikov class, which is the pair
$(\hat{\kappa}_{L},\hat{\kappa}_{R})\in H^{3}(B\text{SU}(N)^{2};\mathrm{U}(1))\cong\mathbb{Z}\times\mathbb{Z}\ .$ (19)
@@ 3 From infrared to ultraviolet: a no-go theorem
One might try to UV complete this ‘product coset’ model with a QCD-like strongly coupled gauge theory, with two types of quark field that confine at different scales to give the two types of pions in the IR. We have already seen that such a UV completion does not fix the mixed WZW term coefficient by anomaly matching, because there is no non-abelian mixed anomaly in the UV. Now we will show something even stronger, which is that the mixed WZW term in the IR is in fact inconsistent with the QCD-like dark sector completion! We will then build a UV completion that does work in §4.
@@
But the absence of a 1-form symmetry means the current algebra for the 0-form QCD symmetries does not close; recall the Postnikov class appearing in the 2-group current algebra relations (28, 29) is integer-quantized, and so preserved under RG flow. Put in a more general context, the problem is that the IR theory (with non-zero mixed WZW term) encodes a non-trivial extension between the flavour symmetry and the 1-form symmetry, manifest in the fibration
@@
One can refine this picture to consider UV theories in which both the 0-form flavour symmetry and the winding number 1-form symmetry are emergent in the infrared. The 2-group emergence theorem [^8] This 2-group emergence theorem has been applied, for instance, to study models for unification in Cordova:2022qtz and, perhaps more closely to the present work, to theories with axions in Brennan:2020ehu ; Choi:2022fgx . of Ref. Cordova:2018cvg then implies the scale hierarchy $\Lambda_{\text{flavour}}\lesssim\Lambda_{\text{1-form}}$, where $\Lambda_{\text{flavour}}$ ($\Lambda_{\text{1-form}}$) denotes the scale where the 0-form QCD flavour symmetry (1-form symmetry) emerges. Assuming the 0-form symmetry remains an exact symmetry of the UV (as in a QCD-like completion with explicit symmetry breaking sources turned off) corresponds to the limit $\Lambda_{\text{flavour}}\to\infty$.
@@ 4.4 Phase structure of QCD coupled to scalar electrodynamics
    $\bm{0<\mu^{2}<1}$. In this case, chiral symmetry breaking occurs in the QCD sector while the scalars remain dynamical, and the photon (from gauging the mediator $\mathrm{U}(1)$) remains massless. The IR theory consists of QCD mesons and baryons, with the baryons coupled weakly via the photon to two complex scalars of mass $m_{\phi}\sim\sqrt{\mu^{2}\Lambda_{\mathrm{QCD}}^{2}}<\Lambda_{\mathrm{QCD}}$. Going to the deep IR, one would also integrate out the baryons and scalars and obtain a free theory of weakly-interacting massless mesons and photons.

### v2

@@ 1 Introduction
Fast forward half a century, and symmetry continues to offer rich and surprising insights into the dynamics of QFTs. A key development in the last decade has been the discovery of generalised notions of symmetry Gaiotto:2014kfa , beyond the action of groups on local operators. One example is higher $p$-form symmetries, in which abelian groups act not on local operators but on extended objects. From the viewpoint of current algebra, the corresponding symmetry currents for $p$-form symmetries (in the case of continuous symmetry) are not vectors $j^{(1)}_{\mu}$ (better, 1-forms) as they are for ‘ordinary’ 00-form symmetries, but tensors $j^{(p+1)}_{\mu...\sigma}$ with more than one index (specifically, $(p+1)$-forms). Returning to QCD and its discrete symmetries, $\text{SU}(n)$ gauge theory possesses such a $\mathbb{Z}_{n}$ valued 1-form symmetry, which was shown by Gaiotto, Kapustin, Komargodski and Seiberg to have a mixed anomaly with parity when the QCD theta angle equals $\pi$ Gaiotto:2017yup , for $n$ even. This subtle anomaly in QCD, discovered nearly 50 years after the ABJ anomaly, means that QCD at $\theta=\pi$ cannot be trivial in the deep infrared (IR).
It has also been recently understood that there can be a non-trivial current algebra between higher-form symmetries of different degree. The simplest example is known as a 2-group symmetry structure, wherein 0-form and 1-form symmetries ‘mix’, as first studied in topological field theory by Kapustin and Thorngren Kapustin:2013uxa , based on the idea of higher gauge theory Baez:2004in ; Baez:2005qu ; Baez:2010ya . This was further elucidated in the wider context of generalised symmetries by Sharpe Sharpe:2015mja . In the case where both 0-form and 1-form symmetries are continuous, this can be captured, following the tradition of current algebra, by a Ward identity of the form Cordova:2020tij :
@@ 2 Mixed WZW term on $\mathrm{SU}(N)\times S^{2}$
Classifications of Wess–Zumino–Witten (WZW) terms Wess:1971yu ; Witten:1983tw via generalized cohomology Freed:2006mx ; Davighi:2018inx ; Lee:2020ojw ; Yonekura:2020upo tell us there is a mixed WZW interaction coming from the existence of a $G$-invariant closed 5-form involving pions on both the $\text{SU}(N)$ and $S^{2}$ factors, namely Davighi:2024zip
@@
where $\mathrm{Vol}_{S^{2}}$ is the volume form on the $S^{2}$. To define this term it is easier to start with a homological description – we briefly discuss the refinement via bordism afterwards. Given there are no homologically non-trivial 4-cycles in the target space, since $H_{4}(\text{SU}(N)\times S^{2})=0$, any 4-cycle in the target space obtained by pushing forward $\Sigma_{4}$ (more precisely, pushing forward a cycle in the fundamental class $[\Sigma_{4}]$) can be realised as the boundary of a 5-cochain $X_{5}$, on which we can integrate $\omega$ to obtain the exponentiated action à la Witten Witten:1983tw
@@
${\cal L}=\frac{in\epsilon^{\mu\nu\rho\sigma}}{48\pi^{2}f_{D}^{2}f_{\pi}^{3}}f_{abc}\epsilon_{ij}\pi_{a}\partial_{\mu}\pi_{b}\partial_{\nu}\pi_{c}\partial_{\rho}\chi_{i}\partial_{\sigma}\chi_{j}+\mathcal{O}(\pi^{4}\chi^{2},\pi^{3}\chi^{3})\,,$ (7)
@@
2.1 Remarks on bordism vs homology, and a discrete theta angle
@@
Roughly speaking, bordism tells us whether manifolds equipped with certain structures — gauge bundles, spin structures, metrics — can be written as boundaries of manifolds in one dimension higher, with all structures smoothly extended thereto. In short, bordism deals directly with manifolds themselves, rather than passing to cochains as in the homological description above. The UV theory we have in mind will feature fermions, so the appropriate bordism theory is (reduced) spin-bordism.
@@
$\tilde{\Omega}_{4}^{\mathrm{Spin}}\left(\text{SU}(N)\times S^{2}\right)\cong\tilde{\Omega}_{4}^{\mathrm{Spin}}\left(S^{2}\right)\cong\mathbb{Z}_{2}\,.$ (8)
This means there is a class of spin 4-manifolds $[X_{4}]$, together with maps $\sigma$ to $\text{SU}(N)\times S^{2}$, that cannot be extended to spin 5-manifolds also equipped with such maps. This equivalence class is the generator of the bordism group (8). The existence of manifolds that are not boundaries is an obstruction to a Witten-like definition (3) of the WZW term when evaluated on such manifolds. This obstruction is not seen using homology.
We can nonetheless proceed in attempting to define our WZW term on a general manifold as follows. That the bordism group is $\mathbb{Z}_{2}$ means that the union of two generators is a boundary of some closed 5-manifold $X_{5}$ with the map $\sigma$ extended. The exponentiated action evaluated on this union $X_{4}\sqcup X_{4}$ must be the square of the action evaluated on the original $X_{4}$, by locality. Therefore, we can define the exponentiated WZW action for this union as
@@
Taking the square root of the right-hand-side then gives two branches of solution, and we must make a consistent choice of branch to define the theory on all possible 4-manifolds $X_{4}$ equipped with background structures. This choice, which is equivalent to specifying a $\mathbb{Z}_{2}$-valued parameter of the theory (also known as a ‘discrete theta angle’), can be made by fixing a representative $X_{4}^{(0)}$ for the non-trivial class in $\tilde{\Omega}_{4}^{\mathrm{Spin}}\left(\text{SU}(N)\times S^{2}\right)\cong\mathbb{Z}_{2}$ and choosing a branch for $\exp(iS[X^{(0)}_{4}])$. For instance, we can fix $X_{4}^{(0)}$ to be a generator of $\tilde{\Omega}_{4}^{\mathrm{Spin}}\left(S^{2}\right)\cong\mathbb{Z}_{2}$ with the background $\text{SU}(N)$ pion configuration set to the trivial map i.e. $\pi_{a}(x)=0$, in which case the RHS of (9) evaluates to unity. Then we can choose either
$\exp(iS[X_{4}^{(0)}])=\pm 1\,.$ (10)
For other manifolds in the same bordism class, for which the topological term in general evaluates to a $\mathbb{C}$-number, the mixed WZW can be defined using a bordism between $X^{(0)}_{4}$ and $X_{4}$, and integrating the differential form $\omega$ thereon.
The choice of sign in (10) is in fact fixed by non-perturbative ’t Hooft anomaly matching: the solution with a minus sign must be chosen if the symmetry group $\text{SU}(2)$ suffers from the mod 2 global anomaly of Witten Witten:1982fp in the UV. To see this, let us be more explicit in our choice of representative and take $X_{4}^{(0)}$ to be $S^{4}$, equipped with a homotopically non-trivial dark pion map $U_{\chi}:S^{4}\to\text{SU}(2)/\mathrm{U}(1)\cong S^{2}$, $[U_{\chi}]=1\mod 2\in\pi_{4}(S^{2})\cong\mathbb{Z}_{2}$. The negative sign choice in (10), which is the non-trivial value of this discrete theta angle, corresponds to assigning a partition function phase $(-1)^{[U_{\chi}]}$ that depends on the homotopy class of the dark pion map, while the positive sign choice assigns the trivial phase to all maps. Now let us discuss how this matches the Witten anomaly. Under a local $\text{SU}(2)$ transformation specified by map $g(x):S^{4}\to\text{SU}(2)\cong S^{3}$, the pion field transforms under the left-action $U_{\chi}\to g\cdot U_{\chi}$. We know that, if the UV theory has non-zero Witten anomaly, then the partition function should flip sign under background gauge transformations for which $[g]=1\mod 2\in\pi_{4}(S^{3})$, for any value of the background gauge field (including zero). Our negative-sign discrete theta angle has this property, as follows. Using the Hopf fibration $f:S^{3}\to S^{2}$, we can pullback $U_{\chi}$ to a map $f^{\ast}U_{\chi}:S^{4}\to S^{3}$, and the associated long exact sequence in homotopy tells us that $[f^{\ast}U_{\chi}]$ is non-trivial in $\pi_{4}(S^{3})$ iff $U_{\chi}$ is non-trivial in $\pi_{4}(S^{2})$. Thus we can write our discrete WZW term as $(-1)^{[f^{\ast}U_{\chi}]}$. Under $g(x)$, this is sent to $(-1)^{[g\cdot(f^{\ast}U_{\chi})]}=(-1)^{[g]\cdot[f^{\ast}U_{\chi}]}=(-1)^{[g]}(-1)^{[f^{\ast}U_{\chi}]}$, thus the partition function flips sign, matching the anomaly. On the other hand for the trivial value of the discrete theta angle, i.e. the positive sign choice, one simply assigns the trivial phase to all paths and so there is no anomalous variation. For more detail on matching global anomalies with discrete WZW terms, we refer the reader to Refs. Yonekura:2020upo ; Lee:2020ojw .
@@ 2.2 An anomaly-matching puzzle
${\cal L}\sim\frac{1}{f_{\pi}f_{D}^{2}}\epsilon^{\mu\nu\rho\sigma}\pi_{0}F_{\mu\nu}\partial_{\rho}\chi_{1}\partial_{\sigma}\chi_{2}\,,$ (11)
@@
${\cal L}\sim\frac{1}{f_{\pi}}\epsilon^{\mu\nu\rho\sigma}\pi_{0}F_{\mu\nu}(F_{D})_{\rho\sigma}\,.$ (12)
@@ 2.3 Infrared 2-group symmetry
$iS[\Sigma_{4}=\partial X_{5}]=\int_{X_{5}}\frac{-in}{8\pi^{2}}\left[\mathrm{CS}(A_{L})-\mathrm{CS}(A_{R})\right]\wedge\mathrm{Vol}_{S^{2}}$ (16)
@@
Mathematically, this (together with 17) defines the infinitesimal version of a continuous 2-group gauge transformation, [^6] We will not actually define what a 2-group is in this paper. For an introduction to the mathematical notion of a 2-group, and how these structures appear in QFT, see e.g. §2 of Davighi:2023luh . with the locally-defined differential forms $(A_{L},A_{R},B)$ forming a 2-connection. Given the 0-form flavour symmetry $\text{SU}(N)^{2}$ and the 1-form winding number symmetry valued in $\mathrm{U}(1)^{[1]}$, possible 2-group structures intertwining the two are classified (subject to certain simplifying assumptions that apply here) by a topological invariant called the Postnikov class, which is the pair
$(\hat{\kappa}_{L},\hat{\kappa}_{R})\in H^{4}(B\text{SU}(N)^{2};\mathbb{Z})\cong\mathbb{Z}\times\mathbb{Z}\ .$ (20)
@@ 3 From infrared to ultraviolet: a no-go theorem
One might try to UV complete this ‘product coset’ model with a QCD-like strongly coupled gauge theory, with two types of quark field that confine at different scales to give the two types of pions in the IR. We have already seen that such a UV completion does not fix the mixed WZW term coefficient by anomaly matching, because there is no non-abelian mixed anomaly in the UV. Now we will show something even stronger, which is that the mixed WZW term in the IR is in fact inconsistent with the QCD-like dark sector completion. We will then build a UV completion that does work in §4.
@@
But the absence of a 1-form symmetry means the current algebra for the 0-form QCD symmetries does not close; recall the Postnikov class appearing in the 2-group current algebra relations (29, 30) is integer-quantized, and so preserved under RG flow. Put in a more general context, the problem is that the IR theory (with non-zero mixed WZW term) encodes a non-trivial extension between the flavour symmetry and the 1-form symmetry, manifest in the fibration
@@
One can refine this picture to consider UV theories in which both the 0-form flavour symmetry and the winding number 1-form symmetry are emergent in the infrared. The 2-group emergence theorem [^8] This 2-group emergence theorem has been applied, for instance, to study models for unification in Cordova:2022qtz and, perhaps more closely to the present work, to theories with axions in Brennan:2020ehu ; Choi:2022fgx . of Ref. Cordova:2018cvg then implies the scale hierarchy $\Lambda_{\text{flavour}}\lesssim\Lambda_{\text{1-form}}$, where $\Lambda_{\text{flavour}}$ ($\Lambda_{\text{1-form}}$) denotes the scale where the 0-form QCD flavour symmetry (1-form symmetry) emerges. Assuming the 0-form symmetry remains an exact symmetry of the UV, as in a QCD-like completion with explicit symmetry breaking sources turned off, corresponds to the limit $\Lambda_{\text{flavour}}\to\infty$.
One might object, however, that the Landau pole associated with the abelian gauge field means that the weakly coupled theory we construct in §4 is not really UV ‘complete’. In §4.5 we sketch how that theory could be further embedded in an asymptotically-free gauge theory, and see how this explicitly entails breaking of the flavour symmetry at finite $\Lambda_{\text{flavour}}$.
@@ 4.4 Phase structure of QCD coupled to scalar electrodynamics
    $\bm{0<\mu^{2}<1}$. In this case, chiral symmetry breaking occurs in the QCD sector while the scalars remain dynamical, and the photon (from gauging the mediator $\mathrm{U}(1)$) remains massless. The IR theory consists of QCD mesons and baryons, with the baryons coupled weakly via the photon to two complex scalars of mass $m_{\phi}\sim\sqrt{\mu^{2}\Lambda_{\mathrm{QCD}}^{2}}<\Lambda_{\mathrm{QCD}}$. Going to the deep IR, one would also integrate out the baryons and scalars and obtain a theory of weakly-interacting massless mesons and photons.
@@
4.5 Deeper into the UV
Because of the Landau pole associated to the abelian gauge field, the short-distance phase we have set out is arguably not a true UV completion. To address this, we here sketch how the abelian gauge theory presented could be further UV completed into a semi-simple gauge theory. But of course, in accordance with the no-go theorem of §3, this necessitates the quark flavour symmetry be only emergent in the abelian gauge theory phase.
The idea will be that the $\text{SU}(n_{c})\times\mathrm{U}(1)$ gauge symmetry emerges from spontaneously breaking a larger $\text{SU}(n_{c}+1)$ gauge symmetry that is linearly realised in the deep UV. Let us embed $\text{SU}(n_{c})$ as the upper left $n_{c}\times n_{c}$ block in the defining representation of $\text{SU}(n_{c}+1)$, and take $\mathrm{U}(1)$ to be generated by the traceless matrix $\mathrm{diag}(1,\dots,1,-n_{c})\subset\mathfrak{su}(n_{c}+1)$. The key consideration will be the fermion sector, its flavour symmetries, and mixed anomalies with $\mathrm{U}(1)$. Recall that we have $N$ flavours of left- and right-handed quark fields $\Psi_{i}$ in the fundamental representation of $\text{SU}(n_{c})$ and with unit $\mathrm{U}(1)$ charge, as well as a pair of complex scalars that are $\text{SU}(n_{c})$ singlets and with unit $\mathrm{U}(1)$ charge also. Now extend this field content by $N$ flavours of left- and right-handed ‘lepton’ fields, that are $\text{SU}(n_{c})$ singlets but carry $\mathrm{U}(1)$ charge $-n_{c}$. The global symmetry of the fermion sector is now enhanced from the $\text{SU}(N)_{L}^{q}\times\text{SU}(N)_{R}^{q}$ quark symmetry of before (now with $q$ for ‘quark’ labels in superscript), that participated in the 2-group structure matched in the deep IR by the mixed WZW term, to a larger
$\text{SU}(N)_{L}^{q}\times\text{SU}(N)_{R}^{q}\times\text{SU}(N)_{L}^{l}\times\text{SU}(N)_{R}^{l}$ (77)
flavour symmetry. If we turn on independent background gauge fields for all four of these $\text{SU}(N)$ factors, the anomaly polynomial contains a piece
$\Phi_{6}=\frac{n_{c}}{2}\frac{h}{2\pi}\left[\mathrm{Tr~{}}\left(\frac{F_{L}^{q}}{2\pi}\right)^{2}-\mathrm{Tr~{}}\left(\frac{F_{R}^{q}}{2\pi}\right)^{2}-\mathrm{Tr~{}}\left(\frac{F_{L}^{l}}{2\pi}\right)^{2}+\mathrm{Tr~{}}\left(\frac{F_{R}^{l}}{2\pi}\right)^{2}\right]\,.$ (78)
Note that the overall coefficient of $n_{c}$ comes from the sum over colour components in the case of the quark piece, while it comes from the ratio of $\mathrm{U}(1)$ charges for the lepton piece. We see that, for the diagonal subgroup of flavour transformations
$G_{\text{diag}}:=\text{SU}(N)_{L}^{q+l}\times\text{SU}(N)_{R}^{q+l}\,,$ (79)
the 2-group coefficients vanish. Accordingly, symmetry matching poses no obstruction to a semi-simple unification pattern (for which there is no 1-form symmetry) whereby quarks and leptons are unified. Sure enough, the $\mathrm{U}(1)$ charge assignment is chosen such that quarks and leptons package into the fundamental representation of $\text{SU}(n_{c}+1)$.
Starting from this deeper UV theory and running down, we have that the separate quark and lepton flavour symmetries emerge at the scale where $\text{SU}(n_{c}+1)$ is Higgsed down to $\text{SU}(n_{c})\times\mathrm{U}(1)$, alongside the magnetic 1-form symmetry associated to the $\mathrm{U}(1)$, and that these 0-form and 1-form global symmetries are fused into 2-group symmetry.
Of course, to write an explicit model the scalars must also be embedded in representations of $\text{SU}(n_{c}+1)$, which necessitates the inclusion of additional coloured scalars, and a mechanism must be put forth for decoupling the extra scalars and the leptons, but we are content to postpone such considerations here.